# GigaBrain-0: A World Model-Powered Vision-Language- Action Model

### v1

@@ Abstract
Training Vision-Language-Action (VLA) models for generalist robots typically requires large-scale real-world robot data, which is expensive and time-consuming to collect. The inefficiency of physical data collection severely limits the scalability, and generalization capacity of current VLA systems. To address this challenge, we introduce GigaBrain-0, a novel VLA foundation model empowered by world model-generated data (e.g., video generation, real2real transfer, human transfer, view transfer, sim2real transfer data). By leveraging world models to generate diverse data at scale, GigaBrain-0 significantly reduces reliance on real robot data while improving cross-task generalization. Our approach further improves policy robustness through RGBD input modeling and embodied Chain-of-Thought (CoT) supervision, enabling the model to reason about spatial geometry, object states, and long-horizon dependencies during task execution. This leads to substantial gains in real-world performance on dexterous, long-horizon, and mobile manipulation tasks. Extensive experiments demonstrate that GigaBrain-0 achieves superior generalization across variations in appearances (e.g., textures, colors), object placements, and camera viewpoints. Additionally, we present GigaBrain-0-Small, an optimized lightweight variant designed to run efficiently on devices such as the NVIDIA Jetson AGX Orin.
@@ 1 Introduction
Recent advances in vision-language-action (VLA) models (Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025a; Bjorck et al., 2025; Cheang et al., 2025; Zhai et al., 2025; Jiang et al., 2025a) have shown promising results in enabling generalist robots to understand high-level instructions, perceive their environments, and execute complex manipulation tasks. These models aim to bridge the gap between symbolic reasoning and embodied action by integrating visual inputs, natural language commands, and motor control into a unified framework. However, training such models typically relies on large-scale datasets of real-world robot interactions, and such data collection is not only expensive and time-consuming but also limited in diversity and scalability. The inefficiency of physical data collection poses a fundamental bottleneck to the development of robust, general-purpose robotic systems capable of operating across a wide range of environments, objects, and task configurations.
@@ 2.1 Vision-Language-Action Models
Developing general-purpose robotic manipulation policies that can interpret high-level instructions and operate effectively in diverse physical environments remains a fundamental challenge in robotics. Prior approaches (Nair et al., 2022; Radford et al., 2021; Xiao et al., 2022; Karamcheti et al., 2023) have focused on leveraging heterogeneous datasets to learn rich, transferable representations that enhance policy robustness in complex and unseen scenarios. Inspired by advances in large language models, VLA frameworks (O’Neill et al., 2024a; Team et al., 2024; Kim et al., 2024; Black et al., 2024; Pertsch et al., 2025; Intelligence et al., 2025; Doshi et al., 2024; Wang et al., 2024a; Liu et al., 2024; Qu et al., 2025; Li et al., 2024; Bjorck et al., 2025; Cheang et al., 2025) have emerged as a promising paradigm, scaling model capacity and data volume to improve generalization across tasks and embodiments. These models typically build upon pretrained vision-language models (Bai et al., 2025; Beyer et al., 2024; Peng et al., 2023; Liu et al., 2023; Alayrac et al., 2022), integrating multimodal inputs to generate action sequences, either through autoregressive token prediction or continuous action modeling via flow matching (Lipman et al., 2022; Liu, 2022). This shift from purely vision-to-action architectures to semantically grounded, language-conditioned policies has significantly expanded the scope of achievable behaviors. To support such data-hungry models, researchers have increasingly relied on cross-embodiment datasets (O’Neill et al., 2024a; Khazatsky et al., 2024; Dasari et al., 2019; Walke et al., 2023; Ebert et al., 2021), combining publicly available logs from heterogeneous robotic platforms as well as large proprietary collections, such as those used in (Cheang et al., 2025; Black et al., 2024; Intelligence et al., 2025; Bjorck et al., 2025). Despite the performance gains, this reliance on extensive real-world interaction data raises practical concerns regarding scalability and cost.
@@ 2.2 World Models as Data Engines
The recent advancement in world model development (Kong et al., 2024; Wang et al., 2025a; Alhaija et al., 2025; Agarwal et al., 2025; Assran et al., 2025; Liao et al., 2025; Jiang et al., 2025b; Jang et al., 2025) has catalyzed a shift toward using synthetic data to narrow the sim-to-real divide in embodied intelligence (Zhu et al., 2024). In fields like autonomous driving, generative models are increasingly employed to simulate complex traffic scenarios, as evidenced by approaches such as (Wang et al., 2024b; Zhao et al., 2025b; Hu et al., 2023; Russell et al., 2025; Gao et al., 2023, 2024; Zhao et al., 2024; Ni et al., 2024; Zhao et al., 2025a; Ren et al., 2025; Ni et al., 2025). In robotics, where data acquisition is often limited by hardware and labor intensity, generative world models offer a promising alternative. Works including (Yang et al., 2023; Du et al., 2023; Zhou et al., 2024; Feng et al., 2025; Jang et al., 2025) leverage natural language prompts to generate plausible future trajectories, which are then converted into low-level control signals via inverse dynamics estimation. To improve the structural fidelity of synthesized outputs, TesserAct (Zhen et al., 2025) and Robot4DGen (Liu et al., 2025b) introduce a unified multimodal generation framework that produces synchronized streams of RGB frames, depth maps, surface normals, or 3D point clouds. This enables the construction of temporally coherent 4D reconstructions, significantly advancing policy learning compared to models trained on RGB video alone. Further enhancements in scene diversity are achieved through techniques like background inpainting (Fang et al., 2025; Yuan et al., 2025), which alters environmental textures, and video-to-video translation methods (Liu et al., 2025a; Wang et al., 2025b; Alhaija et al., 2025; Li et al., 2025; Xu et al., 2025; Dong et al., 2025) that stylize or adapt visual dynamics. Notably, GigaBrain-0 exploits the generative capacity of world models to produce highly varied data across texture, material, illumination, object placement, and camera viewpoints, providing rich, generalizable data source for training VLA models.
@@ 3 GigaBrain-0 Model
GigaBrain-0 is an end-to-end VLA model $g_{\theta}$ that, given visual observations and high-level language instructions, reasons over embodied scenarios to generate compliant action sequences for controlling a wheeled bi-manual robot (e.g., Agilex, G1). To enhance prompt-following fidelity and enable smoother action generation, as illustrated in Fig. 2, GigaBrain-0 adopts a mixture-of-transformers architecture, it leverages a pretrained Vision-Language Model (VLM), PaliGemma2 (Steiner et al., 2024), to encode multimodal inputs, and employs an action Diffusion Transformer (DiT) (Peebles and Xie, 2023) with flow matching (Lipman et al., 2022) to predict action chunks. This hybrid architecture enables decoupled yet synergistic processing of semantic understanding and continuous action generation. During training, we introduce Knowledge Insulation (Driess et al., 2025) to mitigate interference between continuous action-space learning and the VLM’s semantic reasoning capabilities. Furthermore, we augment the VLM head with discrete action token prediction (Pertsch et al., 2025), which significantly accelerates pretraining convergence.
Figure 2: The framework of GigaBrain-0. GigaBrain-0 takes RGB-D input to enhance spatial perception and outputs Embodied Chain-of-Thought (Embodied CoT) as an intermediate representation to strengthen embodied reasoning for manipulation. During training, GigaBrain-0 employs Knowledge Insulation (Driess et al., 2025) to prevent interference between the optimization processes of action prediction and Embodied CoT generation.
To enhance spatial reasoning capabilities, we incorporate RGB-D data during pretraining. Given an input tensor of shape $B\times H\times W\times 4$ (RGB + depth), we first normalize the input and extract visual features using SigLIP (Zhai et al., 2023). To adapt SigLIP to RGB-D inputs, we extend its first convolutional layer with zero-initialized kernels for the depth channel. This preserves pretrained RGB feature extraction while enabling depth-aware representation learning. Notably, SigLIP remains fully trainable throughout GigaBrain-0’s training, allowing adaptive fine-tuning to embodied RGB-D perception. During training, we randomly drop the depth channel (replacing it with zero padding) to ensure compatibility with RGB-only inputs during inference.
Inspired by Chain-of-Thought (CoT) reasoning in LLMs (Wei et al., 2022), we introduce Embodied CoT to improve GigaBrain-0’s reasoning in embodied environments. Unlike standard LLMs, GigaBrain-0 explicitly generates intermediate reasoning tokens, including: (1) manipulation trajectories: 2D projections of end-effector paths onto the image plane, represented by 10 uniformly sampled keypoints; (2) subgoal language: natural language descriptions of intermediate objectives; and (3) discrete action tokens: discrete representations that accelerate training convergence of the subsequent DiT-based continuous action chunk prediction (Pertsch et al., 2025).
@@
where $\mathcal{D}$ denotes the training dataset. $\tau\in[0,1]$ is the flow-matching timestep, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise. $a_{\text{chunk}}^{\tau,\epsilon}=\tau\cdot a_{\text{chunk}}+(1-\tau)\cdot\epsilon$ is the noised action chunk used in flow matching. $M_{\text{CoT},j}\in\{0,1\}$ is a per-token mask indicating whether position $j$ belongs to the CoT reasoning stream (subgoal language or discrete actions). $\hat{\mathbf{t}}_{1:10}$ and $\mathbf{t}_{1:10}$ denote predicted and ground-truth 2D trajectory keypoints, respectively. $\lambda=1$ is a hyperparameter balancing the trajectory regression loss. Notably, we do not manually assign loss weights to the language and action prediction terms, as Knowledge Insulation (Driess et al., 2025) inherently prevents interference between their optimization processes, allowing each stream to learn independently.
@@ 4 GigaBrain-0 Data
For Vision-Language-Action (VLA) models, training data diversity is paramount (Shi et al., 2025). While numerous public datasets (O’Neill et al., 2024b; Wu et al., 2024; Bu et al., 2025b) exist, they are often insufficient in scene variation, or task complexity to enable robust generalization. Real-world robot data collection suffers from high operational costs, low scaling efficiency, and limited environmental diversity, as most deployments repeatedly sample from the same narrow set of scenes.
Recent advances (Bjorck et al., 2025; Jang et al., 2025) have demonstrated that world models can effectively generate diverse and photorealistic training data to augment vision-language-action (VLA) capabilities. In GigaBrain-0, we further extend this paradigm by integrating a broad spectrum of world-model-generated data sources. As shown in Tab. 1, GigaBrain-0 leverages a more diverse set of data sources compared to existing VLA counterparts (Black et al., 2024; Intelligence et al., 2025; Zhai et al., 2025; Jiang et al., 2025a; Cheang et al., 2025; Bjorck et al., 2025). This expanded data diversity significantly reduces reliance on real-world robot-collected data while enhancing model generalization. We detail each data source in the following.
@@ 4.1 Real-World Data
Data Annotation & Processing. For data annotation, if the captured RGB frames lack depth information, we employ MoGe (Wang et al., 2025c) to generate metrically scaled depth maps. Regarding subgoal language annotation, we observe that VLMs struggle to accurately segment long-horizon tasks into meaningful subgoals, while manual segmentation proves prohibitively time-consuming. To address this, we adopt an approach inspired by (James et al., 2020), leveraging gripper state transitions (e.g., open/close, grasp/release) to automatically segment trajectories into atomic subtasks. For each segmented subtask, we utilize Qwen-VL-2.5 (Bai et al., 2025) to generate subgoal language annotations. To mitigate hallucination and ensure consistency, we constrain the annotation process using a structured template and a predefined vocabulary of standardized action phrases (e.g., pick [object], place [object] into [container], open [device]), selected from a curated description library. For 2D manipulation trajectory annotation, we project the 3D end-effector coordinates into the head-mounted camera’s image plane, yielding pixel-space motion traces aligned with visual observations. Notably, we annotate only a subset of the collected data and train our models using a mixture of fully annotated, partially annotated, and raw unannotated trajectories to maximize data utility while managing annotation costs. To further improve pretraining efficiency and reduce redundancy, we perform deduplication across the entire corpus. For each unique task, we retain at most 50 diverse demonstration trajectories, preserving behavioral variety while eliminating near-identical repetitions. This strategy enhances sample efficiency and promotes more robust and generalizable model learning.
@@ 4.2 World-Model-Generated Data
Real2Real Transfer. Real-world robot-collected videos are inherently constrained by the physical environments in which they are captured, including static backgrounds, fixed lighting conditions, and limited variation in object materials, textures, and colors. To overcome these limitations and significantly enhance visual and contextual diversity, we leverage GigaWorld to perform Real2Real Transfer: re-rendering real trajectories in synthetically altered but physically plausible visual contexts. Specifically, we train diffusion-based video generation model (Dong et al., 2025; Liu et al., 2025a) conditioned on geometric and structural priors extracted from real footage. We employ VideoDepthAnything (Chen et al., 2025) to estimate per-frame depth maps and extract Canny edge maps to preserve object boundaries and scene structure. These signals serve as spatial control conditions via a ControlNet (Zhang et al., 2023) branch, enabling precise manipulation of appearance while preserving motion and layout consistency. During inference, for each real-world video clip, we generate approximately 10 visually distinct variants by textually prompting changes to foreground/background materials, surface textures, illumination conditions, and color palettes, all while maintaining the original action semantics and spatial dynamics, as shown in Fig. 4. This approach efficiently multiplies the effective diversity of real data without requiring additional physical collection.
@@
View Transfer. Beyond texture and illumination variation, generalization across observation viewpoints is equally critical for robust embodied perception (Xing et al., 2025). To this end, we exploit GigaWorld’s viewpoint generalization capability to augment real robot-collected data with novel camera perspectives while preserving 3D scene consistency. Specifically, to ensure geometric coherence under viewpoint changes, we project the original RGB frames into novel views using collected depth maps. If depth is unavailable in the source data, we annotate metric-scale depth using MoGe (Wang et al., 2025c). The reprojected views inevitably contain occluded or incomplete regions, we inpaint these using a DiT-based video completion model (Xu et al., 2025) conditioned on the reprojected views. Notably, when the camera viewpoint shifts, the robot’s end-effector must remain functionally consistent with the task, even though its joint configuration changes. We compute the new joint angles via IK based on the updated ego-pose and end-effector pose. The resulting articulated robot geometry is then rendered using its URDF model in a physics-aware simulation engine, and provided as a structural condition (Xu et al., 2025) to the generative model. To mitigate potential discrepancies between simulated and real robot dynamics, we optionally employ a differentiable physics engine (Wang et al., 2025b) to fine-tune motion plausibility and close the sim-to-real physics gap. As shown in Fig. 5, from a single real-world trajectory, our pipeline generates multiple viewpoint-consistent renderings of the scene, complete with dynamically adjusted robot poses that preserve task semantics and physical feasibility.
@@
Sim2Real Transfer. While the methods described above augment real-world data, we further scale our training corpus by leveraging simulation assets to synthesize diverse embodied interaction sequences. Specifically, we compose manipulation scenes in Isaac Sim (NVIDIA, ) using either procedurally generated assets from EmbodiedGen (Wang et al., 2025d) or curated objects from open-source repositories such as ArtVIP (Jin et al., 2025). Robot morphology is defined via URDF files, and end-effector trajectories are computed using IK to ensure physically plausible motion. To bridge the sim-to-real domain gap, particularly in visual appearance, we apply GigaWorld’s Sim2Real Transfer pipeline. Similar in architecture to Real2Real Transfer, this method conditions a diffusion-based video generator (Liu et al., 2025a; Dong et al., 2025) on depth maps exported from the simulation environment. Using text prompts, we dynamically alter surface textures, material reflectance, lighting conditions, and environmental clutter to produce photorealistic renderings that retain the original scene geometry and action semantics, as illustrated in Fig. 6. Crucially, unlike real-world data, simulation affords us full control over scene parameters: we can systematically vary object initial positions, camera viewpoints, background layouts, and even physics properties (e.g., friction, mass) to maximize combinatorial diversity.
@@
Human Video Transfer. Human demonstration videos have emerged as a promising source for training embodied agents (Cheang et al., 2025; Wang et al., 2023; Yang et al., 2025; Kareer et al., 2025; Bu et al., 2025c), offering diversity in tasks, environments, and interaction styles, far exceeding the scale and variety achievable through robotic data collection alone. However, raw human videos exhibit significant gaps when directly applied to robot learning, as egocentric footage often suffers from motion blur, unstable viewpoints, and a visual and kinematic mismatch between human hands and robotic end-effectors. To bridge this gap, we leverage GigaWorld’s video inpainting capabilities to transform large-scale first-person human videos into stable, robot-centric demonstrations. Specifically, we convert videos from the EgoDex dataset (Hoque et al., 2025) into stabilized, robot-executable sequences with articulated mechanical arms replacing human hands. Specifically, we use SAM2 (Ravi et al., 2024) to segment and mask out human hands in each frame. The annotated 3D hand wrist positions (provided in EgoDex) are treated as target end-effector poses for a simulated robot arm. We solve for corresponding joint angles via IK, then render the robot’s URDF model using a physics-aware simulation engine. This rendered arm geometry serves as a structural condition for our diffusion-based generator (Li et al., 2025), ensuring kinematically plausible and visually consistent robot appearances. As shown in Fig. 7, the output is a stabilized, robot-embodied version of the original human demonstration, preserving task intent and spatial relationships while eliminating visual and kinematic domain gaps.
Video Generation with Inverse Dynamics Modeling. Given a single input image, GigaWorld can generate diverse embodied robotic operation videos conditioned on different textual prompts, as illustrated in Fig. 8. Furthermore, we leverage Inverse Dynamics Models (IDM) (Jang et al., 2025) to infer corresponding robot action sequences from these generated videos, which are then used as synthetic training data for embodied manipulation tasks.
@@
Multiview Video Generation. In embodied manipulation scenarios, multiple cameras (e.g., head-mounted and wrist-mounted cameras) are often deployed to capture the operating environment from different viewpoints, creating a demand for generating temporally and spatially consistent multiview videos. To address this, GigaWorld adopts the multiview modeling approach from (Zhao et al., 2025b; Liu et al., 2025a; Dong et al., 2025), which concatenates noise maps from multiple views as input to the diffusion model. This design preserves the original diffusion architecture without modification and enables consistent multiview video generation with only minimal fine-tuning data. As shown in Fig. 9, GigaWorld is capable of generating diverse yet geometrically consistent multiview videos, demonstrating strong cross-view coherence and scene fidelity.

### v3

@@ Abstract
Training Vision-Language-Action (VLA) models for generalist robots typically requires large-scale real-world robot data, which is expensive and time-consuming to collect. The inefficiency of data collection severely limits the scalability, and generalization capacity of current VLA systems. Therefore, we introduce GigaBrain-0, a novel VLA foundation model empowered by world model-generated data. By leveraging world models to generate diverse data at scale, GigaBrain-0 significantly reduces reliance on real robot data while improving cross-task generalization. Our approach further improves policy robustness through RGBD input modeling and embodied Chain-of-Thought (CoT) supervision, enabling the model to reason about spatial geometry, object states, and long-horizon dependencies during task execution. This leads to substantial gains in real-world performance on dexterous, long-horizon, and mobile manipulation tasks. Extensive experiments demonstrate that GigaBrain-0 achieves superior generalization across variations in appearances (e.g., textures, colors), object placements, and camera viewpoints. Additionally, we present GigaBrain-0-Small, an optimized lightweight variant designed to run efficiently on devices such as the NVIDIA Jetson AGX Orin.
@@ 1 Introduction
Recent advances in vision-language-action (VLA) models (pi0; pi05; go1; gr00t; gr3; walloss; galaxea) have shown promising results in enabling generalist robots to understand high-level instructions, perceive their environments, and execute complex manipulation tasks. These models aim to bridge the gap between symbolic reasoning and embodied action by integrating visual inputs, natural language commands, and motor control into a unified framework. However, training such models typically relies on large-scale datasets of real-world robot interactions, and such data collection is not only expensive and time-consuming but also limited in diversity and scalability. The inefficiency of physical data collection poses a fundamental bottleneck to the development of robust, general-purpose robotic systems capable of operating across a wide range of environments, objects, and task configurations.
@@ 2.1 Vision-Language-Action Models
Developing general-purpose robotic manipulation policies that can interpret high-level instructions and operate effectively in diverse physical environments remains a fundamental challenge in robotics. Prior approaches (r3m; radford2021learning; xiao2022masked; karamcheti2023language) have focused on leveraging heterogeneous datasets to learn rich, transferable representations that enhance policy robustness in complex and unseen scenarios. Inspired by advances in large language models, VLA frameworks (o2024open; team2024octo; kimopenvla; pi0; pertsch2025fast; pi05; doshi2024scaling; wang2024scaling; liu2024rdt; qu2025spatialvla; li2024cogact; gr00t; gr3) have emerged as a promising paradigm, scaling model capacity and data volume to improve generalization across tasks and embodiments. These models typically build upon pretrained vision-language models (qwen25vl; paligemma; kosmos; llava; flamingo), integrating multimodal inputs to generate action sequences, either through autoregressive token prediction or continuous action modeling via flow matching (lipman2022flow; liu2022rectified). This shift from purely vision-to-action architectures to semantically grounded, language-conditioned policies has significantly expanded the scope of achievable behaviors. To support such data-hungry models, researchers have increasingly relied on cross-embodiment datasets (o2024open; khazatsky2024droid; dasari2019robonet; walke2023bridgedata; ebert2021bridge), combining publicly available logs from heterogeneous robotic platforms as well as large proprietary collections, such as those used in (gr3; pi0; pi05; gr00t). Despite the performance gains, this reliance on extensive real-world interaction data raises practical concerns regarding scalability and cost.
@@ 2.2 World Models as Data Engines
The recent advancement in world model development (hunyuanvideo; wan; cosmos; cosmospredict; vjepa2; genieEnvisioner; enerverseac; dreamgen) has catalyzed a shift toward using synthetic data to narrow the sim-to-real divide in embodied intelligence (sorasurvey). In fields like autonomous driving, generative models are increasingly employed to simulate complex traffic scenarios, as evidenced by approaches such as (drivedreamer; drivedreamer2; gaia; gaia2; magicdrive; vista; drivedreamer4d; recondreamer; recondreamer++; cosmosdrivedream; recondreamerrl). In robotics, where data acquisition is often limited by hardware and labor intensity, generative world models offer a promising alternative. Works including (unisim; unipi; robodreamer; vidar; dreamgen) leverage natural language prompts to generate plausible future trajectories, which are then converted into low-level control signals via inverse dynamics estimation. To improve the structural fidelity of synthesized outputs, TesserAct (tesseract) and Robot4DGen (Robot4DGen) introduce a unified multimodal generation framework that produces synchronized streams of RGB frames, depth maps, surface normals, or 3D point clouds. This enables the construction of temporally coherent 4D reconstructions, significantly advancing policy learning compared to models trained on RGB video alone. Further enhancements in scene diversity are achieved through techniques like background inpainting (rebot; roboengine), which alters environmental textures, and video-to-video translation methods (robotransfer; embodiedreamer; cosmos; mimicdreamer; egodemogen; emma) that stylize or adapt visual dynamics. Notably, GigaBrain-0 exploits the generative capacity of world models to produce highly varied data across texture, material, illumination, object placement, and camera viewpoints, providing rich, generalizable data source for training VLA models.
@@ 3 GigaBrain-0 Model
GigaBrain-0 is an end-to-end VLA model $g_{\theta}$ that, given visual observations and high-level language instructions, reasons over embodied scenarios to generate compliant action sequences for controlling a wheeled bi-manual robot (e.g., Agilex, G1). To enhance prompt-following fidelity and enable smoother action generation, as illustrated in Fig. 2, GigaBrain-0 adopts a mixture-of-transformers (mot) architecture, it leverages a pretrained Vision-Language Model (VLM), PaliGemma2 (paligemma2), to encode multimodal inputs, and employs an action Diffusion Transformer (DiT) (dit) with flow matching (lipman2022flow) to predict action chunks. This hybrid architecture enables decoupled yet synergistic processing of semantic understanding and continuous action generation. During training, we introduce Knowledge Insulation (KI) to mitigate interference between continuous action-space learning and the VLM’s semantic reasoning capabilities. Furthermore, we augment the VLM head with discrete action token prediction (pertsch2025fast), which significantly accelerates pretraining convergence.
Figure 2: The framework of GigaBrain-0. GigaBrain-0 takes RGB-D input to enhance spatial perception and outputs Embodied Chain-of-Thought (Embodied CoT) as an intermediate representation to strengthen embodied reasoning for manipulation. During training, GigaBrain-0 employs Knowledge Insulation (KI) to prevent interference between the optimization processes of action prediction and Embodied CoT generation.
To enhance spatial reasoning capabilities, we incorporate RGB-D data during pretraining. Given an input tensor of shape $B\times H\times W\times 4$ (RGB + depth), we first normalize the input and extract visual features using SigLIP (siglip). To adapt SigLIP to RGB-D inputs, we extend its first convolutional layer with zero-initialized kernels for the depth channel. This preserves pretrained RGB feature extraction while enabling depth-aware representation learning. Notably, SigLIP remains fully trainable throughout GigaBrain-0’s training, allowing adaptive fine-tuning to embodied RGB-D perception. During training, we randomly drop the depth channel (replacing it with zero padding) to ensure compatibility with RGB-only inputs during inference.
Inspired by Chain-of-Thought (CoT) reasoning in LLMs (cot), we introduce Embodied CoT to improve GigaBrain-0’s reasoning in embodied environments. Unlike standard LLMs, GigaBrain-0 explicitly generates intermediate reasoning tokens, including: (1) manipulation trajectories: 2D projections of end-effector paths onto the image plane, represented by 10 uniformly sampled keypoints; (2) subgoal language: natural language descriptions of intermediate objectives; and (3) discrete action tokens: discrete representations that accelerate training convergence of the subsequent DiT-based continuous action chunk prediction (pertsch2025fast).
@@
where $\mathcal{D}$ denotes the training dataset. $\tau\in[0,1]$ is the flow-matching timestep, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise. $a_{\text{chunk}}^{\tau,\epsilon}=\tau\cdot a_{\text{chunk}}+(1-\tau)\cdot\epsilon$ is the noised action chunk used in flow matching. $M_{\text{CoT},j}\in\{0,1\}$ is a per-token mask indicating whether position $j$ belongs to the CoT reasoning stream (subgoal language or discrete actions). $\hat{\mathbf{t}}_{1:10}$ and $\mathbf{t}_{1:10}$ denote predicted and ground-truth 2D trajectory keypoints, respectively. $\lambda=1$ is a hyperparameter balancing the trajectory regression loss. Notably, we do not manually assign loss weights to the language and action prediction terms, as Knowledge Insulation (KI) inherently prevents interference between their optimization processes, allowing each stream to learn independently.
@@ 4 GigaBrain-0 Data
For Vision-Language-Action (VLA) models, training data diversity is paramount (shi2025diversity). While numerous public datasets (oxe; robomind; agibot) exist, they are often insufficient in scene variation, or task complexity to enable robust generalization. Real-world robot data collection suffers from high operational costs, low scaling efficiency, and limited environmental diversity, as most deployments repeatedly sample from the same narrow set of scenes.
Recent advances (gr00t; dreamgen) have demonstrated that world models can effectively generate diverse and photorealistic training data to augment vision-language-action (VLA) capabilities. In GigaBrain-0, we further extend this paradigm by integrating a broad spectrum of world-model-generated data sources. As shown in Tab. 1, GigaBrain-0 leverages a more diverse set of data sources compared to existing VLA counterparts (pi0; pi05; walloss; galaxea; gr3; gr00t). This expanded data diversity significantly reduces reliance on real-world robot-collected data while enhancing model generalization. We detail each data source in the following.
@@ 4.1 Real-World Data
Data Annotation & Processing. For data annotation, if the captured RGB frames lack depth information, we employ MoGe (moge) to generate metrically scaled depth maps. Regarding subgoal language annotation, we observe that VLMs struggle to accurately segment long-horizon tasks into meaningful subgoals, while manual segmentation proves prohibitively time-consuming. To address this, we adopt an approach inspired by (rlbench), leveraging gripper state transitions (e.g., open/close, grasp/release) to automatically segment trajectories into atomic subtasks. For each segmented subtask, we utilize Qwen-VL-2.5 (qwen25vl) to generate subgoal language annotations. To mitigate hallucination and ensure consistency, we constrain the annotation process using a structured template and a predefined vocabulary of standardized action phrases (e.g., pick [object], place [object] into [container], open [device]), selected from a curated description library. For 2D manipulation trajectory annotation, we project the 3D end-effector coordinates into the head-mounted camera’s image plane, yielding pixel-space motion traces aligned with visual observations. Notably, we annotate only a subset of the collected data and train our models using a mixture of fully annotated, partially annotated, and raw unannotated trajectories to maximize data utility while managing annotation costs. To further improve pretraining efficiency and reduce redundancy, we perform deduplication across the entire corpus. For each unique task, we retain at most 50 diverse demonstration trajectories, preserving behavioral variety while eliminating near-identical repetitions. This strategy enhances sample efficiency and promotes more robust and generalizable model learning.
@@ 4.2 World-Model-Generated Data
Real2Real Transfer. Real-world robot-collected videos are inherently constrained by the physical environments in which they are captured, including static backgrounds, fixed lighting conditions, and limited variation in object materials, textures, and colors. To overcome these limitations and significantly enhance visual and contextual diversity, we leverage GigaWorld to perform Real2Real Transfer: re-rendering real trajectories in synthetically altered but physically plausible visual contexts. Specifically, we train diffusion-based video generation model (emma; robotransfer) conditioned on geometric and structural priors extracted from real footage. We employ VideoDepthAnything (videodepth) to estimate per-frame depth maps and extract Canny edge maps to preserve object boundaries and scene structure. These signals serve as spatial control conditions via a ControlNet (controlnet) branch, enabling precise manipulation of appearance while preserving motion and layout consistency. During inference, for each real-world video clip, we generate approximately 10 visually distinct variants by textually prompting changes to foreground/background materials, surface textures, illumination conditions, and color palettes, all while maintaining the original action semantics and spatial dynamics, as shown in Fig. 4. This approach efficiently multiplies the effective diversity of real data without requiring additional physical collection.
@@
View Transfer. Beyond texture and illumination variation, generalization across observation viewpoints is equally critical for robust embodied perception (xing2025shortcut). To this end, we exploit GigaWorld’s viewpoint generalization capability to augment real robot-collected data with novel camera perspectives while preserving 3D scene consistency. Specifically, to ensure geometric coherence under viewpoint changes, we project the original RGB frames into novel views using collected depth maps. If depth is unavailable in the source data, we annotate metric-scale depth using MoGe (moge). The reprojected views inevitably contain occluded or incomplete regions, we inpaint these using a DiT-based video completion model (egodemogen) conditioned on the reprojected views. Notably, when the camera viewpoint shifts, the robot’s end-effector must remain functionally consistent with the task, even though its joint configuration changes. We compute the new joint angles via IK based on the updated ego-pose and end-effector pose. The resulting articulated robot geometry is then rendered using its URDF model in a physics-aware simulation engine, and provided as a structural condition (egodemogen) to the generative model. To mitigate potential discrepancies between simulated and real robot dynamics, we optionally employ a differentiable physics engine (embodiedreamer) to fine-tune motion plausibility and close the sim-to-real physics gap. As shown in Fig. 5, from a single real-world trajectory, our pipeline generates multiple viewpoint-consistent renderings of the scene, complete with dynamically adjusted robot poses that preserve task semantics and physical feasibility.
@@
Sim2Real Transfer. While the methods described above augment real-world data, we further scale our training corpus by leveraging simulation assets to synthesize diverse embodied interaction sequences. Specifically, we compose manipulation scenes in Isaac Sim (issacsim) using either procedurally generated assets from EmbodiedGen (embodiedgen) or curated objects from open-source repositories such as ArtVIP (artvip). Robot morphology is defined via URDF files, and end-effector trajectories are computed using IK to ensure physically plausible motion. To bridge the sim-to-real domain gap, particularly in visual appearance, we apply GigaWorld’s Sim2Real Transfer pipeline. Similar in architecture to Real2Real Transfer, this method conditions a diffusion-based video generator (robotransfer; emma) on depth maps exported from the simulation environment. Using text prompts, we dynamically alter surface textures, material reflectance, lighting conditions, and environmental clutter to produce photorealistic renderings that retain the original scene geometry and action semantics, as illustrated in Fig. 6. Crucially, unlike real-world data, simulation affords us full control over scene parameters: we can systematically vary object initial positions, camera viewpoints, background layouts, and even physics properties (e.g., friction, mass) to maximize combinatorial diversity.
@@
Human Video Transfer. Human demonstration videos have emerged as a promising source for training embodied agents (gr3; mimicplay; egovla; egomimic; bu2025univla), offering diversity in tasks, environments, and interaction styles, far exceeding the scale and variety achievable through robotic data collection alone. However, raw human videos exhibit significant gaps when directly applied to robot learning, as egocentric footage often suffers from motion blur, unstable viewpoints, and a visual and kinematic mismatch between human hands and robotic end-effectors. To bridge this gap, we leverage GigaWorld’s video inpainting capabilities to transform large-scale first-person human videos into stable, robot-centric demonstrations. Specifically, we convert videos from the EgoDex dataset (egodex) into stabilized, robot-executable sequences with articulated mechanical arms replacing human hands. Specifically, we use SAM2 (sam2) to segment and mask out human hands in each frame. The annotated 3D hand wrist positions (provided in EgoDex) are treated as target end-effector poses for a simulated robot arm. We solve for corresponding joint angles via IK, then render the robot’s URDF model using a physics-aware simulation engine. This rendered arm geometry serves as a structural condition for our diffusion-based generator (mimicdreamer), ensuring kinematically plausible and visually consistent robot appearances. As shown in Fig. 7, the output is a stabilized, robot-embodied version of the original human demonstration, preserving task intent and spatial relationships while eliminating visual and kinematic domain gaps.
Video Generation with Inverse Dynamics Modeling. Given a single input image, GigaWorld can generate diverse embodied robotic operation videos conditioned on different textual prompts, as illustrated in Fig. 8. Furthermore, we leverage Inverse Dynamics Models (IDM) (dreamgen) to infer corresponding robot action sequences from these generated videos, which are then used as synthetic training data for embodied manipulation tasks.
@@
Multiview Video Generation. In embodied manipulation scenarios, multiple cameras (e.g., head-mounted and wrist-mounted cameras) are often deployed to capture the operating environment from different viewpoints, creating a demand for generating temporally and spatially consistent multiview videos. To address this, GigaWorld adopts the multiview modeling approach from (drivedreamer2; robotransfer; emma), which concatenates noise maps from multiple views as input to the diffusion model. This design preserves the original diffusion architecture without modification and enables consistent multiview video generation with only minimal fine-tuning data. As shown in Fig. 9, GigaWorld is capable of generating diverse yet geometrically consistent multiview videos, demonstrating strong cross-view coherence and scene fidelity.